OpenStudy (anonymous):

solve the nonlinear system of equation 3y^2-5x^2=7 and 6y^2-7x^2=26

OpenStudy (turingtest):

\[ 3y^2-5x^2=7\]\[6y^2-7x^2=26\]multiply the first equation by -2 and add it to the second, then solve for x. use what you get for x to find y

jhonyy9 (jhonyy9):

3y^2 -5x^2 =7 6y^2 -7x^2 =26 ---- 6y^2 =7x^2 +26 ---- y^2 =(7x^2 +26)/6 3(7x^2 +26)/6 -5x^2 =7 (7x^2 +26)/2 -5x^2 =7 7x^2 +26 -10x^2 =14 -3x^2 =-12 3x^2 =12 x^2 = 12/3 =4 x_1,2 = +/- sqrt4 = +/- 2 y^2 =(7*4^2 +26)/6 =(7*16 +26)/6 =(112 +26)/6 = 138/6 =23 y_1,2=+/- sqrt23 x_1,2= +/- 2 y_1,2= +/- sqrt23